OpenStudy (anonymous):

Solve 3√2y - 3 = 3√2 - y for y? I will give out a gold metal. I just need the answer. Thank you!

OpenStudy (anonymous):

Is it \[3*\sqrt{2y-3}=3*\sqrt{2-y}\]? if so then the answer is \[y=\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

You are such a life saver!! Thank you!!